OpenStudy (anonymous):

The dimensions of a solar panel are 3.4 feet wide and 6.2 feet long. Match the description on the left with the appropriate calculation. Length WIdth Perimeter Area No Match A. 3.4 ft B. 19.2 ft C. 21.08 ft D. 21.08 sq ft E. 6.2 ft

OpenStudy (anonymous):

@bharad

OpenStudy (raffle_snaffle):

So the length is the longest side while the width is the shorter side

OpenStudy (raffle_snaffle):

what is the formula for parameter of a rectangle?

OpenStudy (anonymous):

so width is 3.4 ft ?

OpenStudy (anonymous):

length= E (longest side) breadth=A(shorter side) perimeter=B(perimeter=2(L+B) ) area=D(L*B)(unit is sq ft) no match = C